I worked a bit on recovering the DRMI locking again tonight. I decided to shutter the AUX laser on the PSL table at least until I figured out the correct locking settings. As has become customary now, there was a cable in the AS beampath (leading from the AS55 DC monitor to nothing, through the enclosure side panel, it is visible in Attachment #3 in this elog) which I only found after 30mins of futility - please try and remove all un-necessary cables and leave the AS beampath in a usable state after working on the AS table! In the end, I got several short (~3mins) stretches in tonight, but never long enough to do the loop characterization I wanted to get in tonight, probably wrong gains in one or more of the loops. In the last 30 minutes, the IMC has been frequently losing lock, so I am quitting for now. The AUX laser remains shuttered.